### v1

# Scalable unsupervised feature selection via weight stability

Abstract
Unsupervised feature selection is critical for improving clustering performance in high-dimensional data, where irrelevant features can obscure meaningful structure. In this work, we introduce the Minkowski weighted $k$-means++, a novel initialisation strategy for the Minkowski Weighted $k$-means. Our initialisation selects centroids probabilistically using feature relevance estimates derived from the data itself. Building on this, we propose two new feature selection algorithms, FS-MWK++, which aggregates feature weights across a range of Minkowski exponents to identify stable and informative features, and SFS-MWK++, a scalable variant based on subsampling. We support our approach with a theoretical guarantee under mild assumptions and extensive experiments showing that our methods consistently outperform existing alternatives.
Keywords: Unsupervised feature selection, clustering, noisy data.
1 Introduction
Clustering is a fundamental machine learning technique that assigns data points in a dataset to clusters, with each cluster containing similar data points. As clustering algorithms operate without the need for labelled data, they have been successfully applied across various domains, including data pre-processing, quantitative finance, image analysis, and bioinformatics [1, 2, 3, 4]. However, their performance (and that of any other machine learning algorithm) can be heavily influenced by the quality and dimensionality of the data. High-dimensional datasets often contain redundant or irrelevant features, which can obscure meaningful patterns and degrade both efficiency and accuracy. Feature selection algorithms mitigate this issue by identifying and retaining only the most informative features [5, 6].
Unfortunately, most existing feature selection algorithms rely on labelled data, which may be noisy, corrupted, or entirely unavailable in many real-world scenarios. This limitation highlights the need for effective unsupervised feature selection methods, which can improve learning performance without requiring labels. Despite growing interest in unsupervised feature selection, scalable methods that perform well across diverse data geometries remain scarce. This is particularly limiting in domains such as genomics, anomaly detection, or remote sensing, where data is high-dimensional and labels are either unavailable or unreliable.
Clustering algorithms can be broadly categorised into different types, including density-based, hierarchical, fuzzy, and partitional methods (see, for instance [7], and references therein). Density-based approaches, such as the classic DBSCAN [8], group points based on density regions, making them effective for arbitrary-shaped clusters. Hierarchical algorithms build a nested hierarchy of clusters, which can be visualised with a dendrogram. Fuzzy clustering algorithms assign each data point to every cluster with different degrees of membership, usually adding to one. Here, we focus on partitional clustering.
Let $X$ be a dataset, where each $x_{i}\in X$ is described over $m$ features. Partitional algorithms produce a clustering $S=\{S_{1},\ldots,S_{k}\}$ such that $X=\bigcup_{l=1}^{k}S_{l}$ and $S_{l}\cap S_{t}=\emptyset$ for $l,t=1,\ldots,k$ with $l\neq t$. $k$-means [9] is, arguably, the most popular partitional clustering algorithm [10, 11]. Like most clustering algorithms, $k$-means assumes that all features contribute equally to cluster formation. However, in many real-world datasets, certain features are more informative than others, and treating them equally can lead to poor clustering performance.
Feature-weighted clustering methods address this limitation by assigning importance weights to features based on their contribution to the underlying clustering structure. Minkowski Weighted $k$-means (MWK) [12] extends $k$-means by incorporating a feature weighting mechanism that adjusts the Minkowski distance metric based on the within-cluster dispersion of features. Although MWK improves cluster recovery, the computed feature weights are sensitive to the Minkowski exponent and initial centroids used (for details, see Section 2). Our work addresses both of these issues by introducing a probabilistic, relevance-aware initialisation and using the stability of feature weights across exponents to guide unsupervised feature selection. In addition to empirical validation, we also provide a theoretical guarantee supporting our proposed approach.
Our main contribution in this paper is threefold. First, we develop a probabilistic initialisation strategy for MWK, guided by feature relevance, that consistently improves clustering performance. Second, we demonstrate that by analysing the stability of feature weights computed under different parameter settings, these weights can be effectively used for unsupervised feature selection. Third, we design a sampling-based variant that enables our method to scale to large, high-dimensional datasets, achieving superior performance across benchmark tasks.
2 Related work
In this section, we describe the related work pertinent to this paper, focusing on two main areas. Section 2.1 presents a general overview of prominent clustering algorithms, while Section 2.2 describes key unsupervised feature selection methods and associated challenges.
2.1 Clustering
The $k$-means algorithm [9] is a widely used partitional clustering method, which has been extended in numerous ways [13, 11]. It produces a partition $S=\{S_{1},\ldots,S_{k}\}$ of a dataset $X$ such that $S_{l}\cap S_{t}=\emptyset$ for all $l,t=1,\ldots k$ with $l\neq t$. This is achieved by minimising the objective function
$W(S,Z)=\sum_{l=1}^{k}\sum_{x_{i}\in S_{l}}\sum_{v=1}^{m}(x_{iv}-z_{lv})^{2},$ (1)
where $z_{l}\in Z$ is the centroid of cluster $S_{l}$. The minimisation follows three straightforward steps: (i) select $k$ data points from $X$ uniformly at random, and assign their values to $z_{1},...,z_{k}$; (ii) assign each $x_{i}\in X$ to the cluster $S_{l}$ whose centroid $z_{l}$ is the nearest to $x_{i}$; (iii) update each $z_{l}\in Z$ to the component-wise mean of the points in $S_{l}$. If there are changes in $Z$, go back to step (ii).
Although popular, $k$-means is not without drawbacks. For instance, it assumes that all features of $X$ are equally relevant. This is problematic because it may lead to an irrelevant feature contributing to the clustering just as much as a relevant feature. Moreover, even among relevant features there may be different degrees of relevance, and a robust algorithm should take this into account. Also, $k$-means is a greedy algorithm that does not guarantee convergence to a global minimum. As a result, the choice of initial centroids is particularly important. Poor initialisation may lead to conversion at a suboptimal solution.
The $k$-means++ [14] algorithm addresses the latter of the problems above. It improves the initial centroids of $k$-means by spreading them out more strategically based on the data distribution. This often reduces the likelihood of poor local minima, leading to better clustering results. The $k$-means++ algorithm is now the default $k$-means implementation in many popular software packages such as MATLAB, scikit-learn, and R. Let $d(x_{i})$ be the distance from $x_{i}$ to its nearest centroid in $Z$. That is,
$d(x_{i})=\min_{z_{l}\in Z}\sum_{v=1}^{m}(x_{iv}-z_{lv})^{2},$
$k$-means++ works by: (i) select a data point $x_{t}\in X$ uniformly at random and set $Z=\{x_{t}\}$; (ii) select a new point $x_{t}\in X$ with probability $\frac{d(x_{t})^{2}}{\sum_{i=1}^{n}d(x_{i})^{2}}$, and set $Z=Z\cup\{x_{t}\}$; (iii) repeat step (ii) until $|Z|=k$.
The $k$-means and $k$-means++ algorithms rely on the Euclidean distance, making them biased towards Gaussian (spherical) clusters [15] and imposing the assumption that all features are equally relevant. The Minkowski weighted $k$-means(MWK) [12] is a popular method (see for instance [16, 17, 18, 19]) that overcomes these issues by introducing cluster-specific feature weights into the Minkowski distance. That is, the distance between a data point $x_{i}$ and a centroid $z_{l}$ is given by
$d_{p}(x_{i},z_{l})={\sum_{v=1}^{m}}w_{lv}^{p}\lvert x_{iv}-z_{lv}\lvert^{p},$
where $p$ is the Minkowski exponent. This leads to the objective function
$W_{p}(S,Z,w)=\sum_{l=1}^{k}\sum_{x_{i}\in S_{l}}\sum_{v=1}^{m}w_{lv}^{p}{\lvert x_{iv}-z_{lv}\rvert^{p}}.$ (2)
To minimise the above, MWK defines the within-cluster dispersion of each feature as $D_{lv}=\sum_{x_{i}\in S_{l}}|x_{iv}-z_{lv}|^{p}$. By re-writing (2) as $\sum_{l=1}^{k}\sum_{v=1}^{m}w_{lv}^{p}D_{lv}$ and minimising it subject to $\sum_{v=1}^{m}w_{lv}=1$ for $l=1,\ldots,k$, the following optimal weights are obtained
$w_{lv}=\frac{1}{\sum_{u=1}^{m}\left[\frac{D_{lv}}{D_{lu}}\right]^{\frac{1}{p-1}}}.$ (3)
In the above, $w_{lv}$ represents the weight of feature $v$ at cluster $S_{l}$. For $p>1$, this weight will be higher in features with values concentrated around the centroid (i.e., those with lower dispersion), thereby reflecting their degree of relevance. This aligns with the intuitive notion that a feature may have different degrees of importance at different clusters. Moreover, employing the Minkowski distance allows MWK to adapt to different cluster shapes, reducing the bias toward Gaussian clusters.
2.2 Unsupervised feature selection
High-dimensional datasets often contain redundant or irrelevant features, which can obscure meaningful patterns, increase computational cost, and impair model generalisation. Feature selection addresses these challenges by identifying such features, thereby leading to more efficient, stable, and reliable results.
Feature selection using feature similarity (FSFS) [20] is one of the most popular unsupervised feature selection methods (as indicated by its high number of citations in Google Scholar). Hence, we include it in our comparison (see Section 5). FSFS aims at identifying and removing redundant features by calculating their maximum information compression index,
$2\lambda_{2}(v_{i},v_{j})=$ $\mathrm{var}(v_{i})+\mathrm{var}(v_{j})$
$-\sqrt{(\mathrm{var}(v_{i})+\mathrm{var}(v_{j}))^{2}-4\,\mathrm{var}(v_{i})\,\mathrm{var}(v_{j})(1-\rho(v_{i},v_{j})^{2})}$ (4)
where $\rho(v_{i},v_{j})$ is the Pearson correlation coefficient between features $v_{i}$ and $v_{j}$. Notice that $\lambda_{2}$ measures the minimum information loss when projecting features to a lower dimension, hence, its use as a measure of information redundancy. FSFS proceeds as follows.
  1. 1.
    Initialize the full feature set $V=\{1,\ldots,m\}$, and choose a parameter $\kappa$ such that $\kappa\leq m-1$.
  2. 2.
    Select the feature $v^{*}\in V$ with the lowest redundancy score $r^{\kappa}_{v}$. Retain $v^{*}$ and remove the $\kappa$ most similar features to $v^{*}$ from $V$.
  3. 3.
    Set $\epsilon=r^{\kappa}_{v^{*}}$, and $\kappa=\min(\kappa,|V|-1)$. If $\kappa=1$, go to Step 6.
  4. 4.
    While $r^{k}_{v}>\epsilon$, do the following:
    1. (a)
      Set $\kappa=\kappa-1$.
    2. (b)
      Set $r^{\kappa}_{v}=\min\limits_{v\in V}r^{\kappa}_{v}$
    3. (c)
      If $\kappa=1$, go to Step 6.
  5. 5.
    Go to Step 2.
  6. 6.
    Return $V$.
Notice that in FSFS, the parameter $\kappa$ refers to the number of nearest neighbours considered. That is, $r^{\kappa}_{v}$ is the dissimilarity between feature $v$ and its $\kappa^{th}$ nearest neighbour. Here, this use of $\kappa$ is unrelated to the $k$ of $k$-means.
Multi-Cluster Feature Selection (MCFS) [21] is another popular unsupervised feature selection method. MCFS combines spectral techniques from manifold learning with $\ell_{1}$-regularization to identify the most relevant features. The algorithm has a single parameter $\kappa$, which specifies the number of neighbors in the $\kappa$-nearest neighbours graph. The original paper suggests setting $\kappa=5$ as a default value. In our experiments (see Section 5), we tuned this parameter from 1 to 5, and report the best results. While this procedure introduces a slight positive bias (i.e., it favours MCFS), it does not compromise our objectives. The algorithm is as follows.
  1. 1.
    Construct a $\kappa$-nearest neighbours graph to model the local manifold structure of the data.
  2. 2.
    Solve a generalised eigen-problem and retain the top $k$ eigenvectors, where $k$ is the number of clusters.
  3. 3.
    Solve $k$ $\ell_{1}$-regularised regression problems using Least Angle Regression (LAR) and obtain $k$ sparse coefficient vectors.
  4. 4.
    Compute an MCFS score for each feature $v=1,\ldots,m$.
  5. 5.
    Return the top features with the highest MCFS scores.
We direct readers interested in further details to the original publication [21].
Subspace Clustering Feature Selection (SCFS) [22] is a recent unsupervised feature selection method combining subspace clustering with sparse regression. SCFS aims to select features that preserve the multi-cluster structure of the data by adaptively learning sample similarities. First, it computes a similarity matrix implicitly by solving a self-expressive model, in which each sample is represented through a low-dimensional space shared with other similar samples. Then, a regularised regression model is applied to link features to the discovered clustering structure. Features with stronger associations to the cluster are ranked higher. SCFS introduces two regularisation paramenters, $\alpha$ and $\beta$, to control the trade-off between similarity preservation and sparsity. Following the original paper we tuned these parameters using grid search over the set $\{10^{-4},10^{-2},1,10^{2},10^{4}\}$ and selected the best performing combination. We direct readers interested in further details about SCFS to its original publication [22].
Laplacian Score-regularized Concrete Autoencoder (LS-CAE) [23] is a recent deep learning-based approach to unsupervised feature selection. LS-CAE extends the Concrete Autoencoder (CAE) framework by incorporating a Laplacian score regularisation term into the loss function. This encourages the selection of features that both enable reconstruction of the original data and preserve the local clustering structure. To achieve this, LS-CAE trains a fully differentiable autoencoder, where feature selection is implemented through a Concrete layer that softly samples features in a learnable manner. The final selected features are obtained after annealing the sampling temperature towards discrete choices.
Although LS-CAE operates without class labels, and is therefore unsupervised, it differs from traditional unsupervised feature selection methods such as FSFS, MCFS, or SCFS. In particular, LS-CAE involves training a deep neural network optimised with multiple objectives, which arguably provides greater modelling capacity compared to classical methods based solely on feature scoring or graph-based analysis. As a result, LS-CAE may potentially have an advantage in certain settings. However, our experimental results (see Section 5) demonstrate that other non-deep-learning methods (such as the one we introduce) can outperform LS-CAE, suggesting that deep architectures are not universally superior for unsupervised feature selection.
3 Our proposed methods
This section is divided into two parts. Section 3.1 introduces a novel initialisation method for the Minkowski Weighted $k$-means (MWK) algorithm. This method selects initial centroids by taking into account the relative relevance of features, thereby improving cluster recovery. Section 3.2 builds upon this foundation to develop a new unsupervised feature selection algorithm. This method analyses the stability of feature weights generated under varying parameter settings. Additionally, we propose a sampling-based extension that allows our method to scale effectively to high-dimensional data.
3.1 The Minkowski weighted $k$-means++
Here, we propose a novel initialisation strategy for the MWK, taking inspiration from $k$-means++ (for details, see Section 2.1). Our method, which we refer as Minkowski Weighted $k$-means++ (MWK++) selects initial centroids by incorporating the relevance of features as estimated via their dispersions.
Algorithm 1 Minkowski Weighted $k$-means++ (MWK++)
1:Dataset $X$, number of clusters $k$, exponent $p>1$.
2:Initial centroids $Z=\{z_{1},\ldots,z_{k}\}$, initial feature weights $w$.
3:Select the first centroid $z_{1}\in X$ uniformly at random, and set $Z=\{z_{1}\}$.
4:Compute the Minkowski centre $c\in\mathbb{R}^{m}$ of the dataset using exponent $p$
5:For each feature $v=1,\ldots,m$ compute the dispersion
$D_{v}=\sum_{i=1}^{n}|x_{iv}-c_{v}|^{p}.$
6:Increment each $D_{v}$ by the average of $D$.
7:Compute feature weights
$w_{v}=\left(\sum_{u=1}^{m}\left(\frac{D_{v}}{D_{u}}\right)^{1/(p-1)}\right)^{-1}.$
8:Replicate $w$ to form a $k\times m$ matrix.
9:for $i=2$ to $k$ do
10:    For each $x_{i}\in X$, compute distance to nearest centroid
$d(x_{i})=\min_{l=1,\ldots,k}\sum_{v=1}^{m}w_{lv}^{p}|x_{iv}-z_{lv}|^{p}.$
11:    Set sampling probability
$P(x_{i})=\frac{d(x_{i})}{\sum_{j=1}^{n}d(x_{j})}.$
12:    Select one $x_{t}\in X$ according to $P(x_{t})$, and add it to $Z$.
13:end for
14:return $Z$, $w$
Our approach improves centroid selection by biasing the sampling process towards regions of the feature space where features have a lower dispersion, and are then more likely to be informative. By computing feature weights prior to sampling, MWK++ ensures that distances are measured with emphasis on relevant features. As a result, the selected centroids tend to be well-separated along the more informative dimensions, increasing the likelihood of high-quality clustering outcomes (for details, see Section 5). In the next section, we show how feature weights computed with MWK++ can also serve as a foundation for our new unsupervised feature selection method.
3.2 Feature Selection With MWK++
This section introduces our Feature Selection method with MWK++ (FS-MWK++). This is a novel unsupervised feature selection method based on the stability of feature weights computed by the MWK++ algorithm across a range of Minkowski exponents. Recall that the Minkowski distance induces different geometric biases depending on the value of $p$. For example, in two dimensions, lower values such as $p=1.1$ tend to favour diamond-shaped clusters, whereas $p=2$ corresponds to the standard Euclidean distance and favours spherical clusters. By evaluating a broad range of $p$ values, our method captures clustering structures under diverse distance biases, thereby reducing the risk of selecting features that are overly tailored to a single geometric bias.
Rather than attempting to optimise $p$ directly, we exploit the variability in clustering outcomes to assess the stability of feature relevance across multiple settings. For each $p$, we run MWK++ multiple times and retain the feature weights corresponding to the lowest objective value. These weight vectors are then aggregated by taking the component-wise median, yielding a robust estimate of each feature’s overall importance (for details, see Algorithm 2). Features that consistently receive high weights across different clustering geometries are more likely to be genuinely informative. This stability-based approach naturally filters out noisy or unstable features without requiring supervision or parameter tuning.
Algorithm 2 Feature Selection with MWK++ (FS-MWK++)
1:Dataset $X$, number of clusters $k$, number of features to select $r$.
2:A subset containing the $r$ most informative features.
3:Define a set of Minkowski exponents $P=\{1.1,\ldots,3.0\}$.
4:for all $p\in P$ do
5:    Run MWK++ 25 times on $X$, independently, with exponent $p$.
6:    From these, identify the clustering that minimises  (2).
7:    Retain the feature weights of the clustering identified in the previous step.
8:end for
9:Compute the component-wise median of the retained feature weights.
10:return The indices of the $r$ features with the highest median weights.
We now formally justify our use of weight stability across different values of $p$ as the basis for FS-MWK++ by presenting a theoretical guarantee. We begin with a short definition and a lemma.
Definition 1.
Let $D_{lv}^{(p)}$ denote the within-cluster dispersion of feature $v$ in cluster $S_{l}$ computed with Minkowski exponent $p$, and let $w_{lv}^{(p)}$ denote the corresponding feature weight.
Lemma 1.
Let $v$ be a noise feature, and suppose that all such features are drawn independently from the same distribution and are uncorrelated with cluster structure. If there exists at least one relevant feature, then for any $p>1$,
$w_{lv}^{(p)}<\frac{1}{m}.$
Proof.
Suppose $u^{*}$ is a relevant feature. Since relevant features are aligned with cluster structure, we have that
$D_{lu^{*}}^{(p)}<D_{lv}^{(p)}.$
Let $r=\left(\frac{D_{lv}^{(p)}}{D_{lu^{*}}^{(p)}}\right)^{\frac{1}{p-1}}>1$. Now consider the sum in the denominator of Equation (3),
$\sum_{u=1}^{m}\left(\frac{D_{lv}^{(p)}}{D_{lu}^{(p)}}\right)^{\frac{1}{p-1}}=r+\sum_{u\neq u^{*}}\left(\frac{D_{lv}^{(p)}}{D_{lu}^{(p)}}\right)^{\frac{1}{p-1}}.$
Since all noise features have dispersions similar to $D_{lv}^{(p)}$, for each irrelevant $u$,
$\left(\frac{D_{lv}^{(p)}}{D_{lu}^{(p)}}\right)^{\frac{1}{p-1}}\approx 1\Rightarrow\sum_{u=1}^{m}\left(\frac{D_{lv}^{(p)}}{D_{lu}^{(p)}}\right)^{\frac{1}{p-1}}\approx(m-1)+r>m\Rightarrow w_{lv}^{(p)}<\frac{1}{m}.$
This lemma allows us to show that, under mild assumptions, relevant features have a higher median weight than noise features when aggregated across clusters and distance exponents. The next theorem formalises this insight.
Theorem 1.
Let $P\neq\emptyset$ be a finite set of valid Minkowski exponents. If $u^{*}$ is a relevant feature for at least half of the clusters and $v$ is a noise feature drawn independently from a common distribution and uncorrelated with cluster structure, then
$\mathrm{median}\left(\{w_{lu^{*}}^{(p)}:p\in P,\,l=1,\dots,k\}\right)>\mathrm{median}\left(\{w_{lv}^{(p)}:p\in P,\,l=1,\dots,k\}\right).$
Proof.
From Lemma 1, we have that for any $p>1$
$w_{lv}^{(p)}\in\{w_{lv}^{(p)}:p\in P,\,l=1,\dots,k\}\Rightarrow w_{lv}^{(p)}<\frac{1}{m}.$
Hence, $\mathrm{median}\left(\{w_{lv}^{(p)}:p\in P,\,l=1,\dots,k\}\right)<\frac{1}{m}$.
Now consider the set $\{w_{lu^{*}}^{(p)}:p\in P,\,l=1,\dots,k\}$. Observe that Equation (3) is strictly convex for $p>1$, so
$w_{lu^{*}}^{(p)}=\frac{1}{m}\iff\sum_{u=1}^{m}\left[\frac{D_{lu^{*}}}{D_{lu}}\right]^{\frac{1}{p-1}}=m\iff\frac{D_{lu^{*}}}{D_{lu}}=1\text{ for all }u=1,\ldots,m.$
However, the existence of $v$ shows $X$ has at least one noise feature. Thus, the above cannot happen and by consequence $w_{lu^{*}}^{(p)}\neq\frac{1}{m}$. Recall that (3) ensures the weights for a particular cluster $S_{l}$ must sum to one, and that $w_{lv}^{(p)}<\frac{1}{m}$. Hence, the difference $\frac{1}{m}-w_{lv}^{(p)}>0$ is spread over any relevant feature, and $w_{lu^{*}}^{(p)}>\frac{1}{m}$. We have, by hypothesis, that $u^{*}$ is relevant for at least half of the clusters. Therefore, $\mathrm{median}\left(\{w_{lu^{*}}^{(p)}:p\in P,\,l=1,\dots,k\}\right)>\frac{1}{m}$. ∎
While FS-MWK++ is effective at identifying stable and informative features (for details, see Section 5), it becomes computationally expensive on large datasets. We address this limitation in three ways. First, we modify the computation of the Minkowski centre. Instead of solving an optimisation problem, we approximate the centre by using the component-wise median when $p<1.5$, and the mean otherwise. This change substantially reduces runtime while preserving good results. Second, we reduce the number of $p$ values in our experiments by setting $P=\{1.1,1.3,1.5,\ldots,3.0\}$. That is, we have a set of 10 equally spaced values containing 1.1 and 3.0. Third, we introduce a scalable variant that further reduces the computational cost by operating on a representative subset of the data. This sampling-based alternative retains the core idea of stability across exponents while significantly improving runtime efficiency. Algorithm 3 describes the steps of this method.
Algorithm 3 Sample Feature Selection with MWK++ (SFS-MWK++)
1:Dataset $X\in\mathbb{R}^{nm}$, number of clusters $k$, number of features $r$.
2:A subset containing the $r$ most informative features.
3:Define a set of Minkowski exponents $P=\{1.1,1.3,1.5,\ldots,3.0\}$.
4:Set the sample size $n_{s}=k\sqrt{n}$.
5:for i=1 to 25 do
6:    Create a dataset $X_{s}$ containing $n_{s}$ data points from $X$ selected uniformly at random.
7:    for all $p\in P$ do
8:         Run MWK++ 25 times on $X$, independently, with exponent $p$.
9:         From these, identify the clustering that minimises  (2).
10:         Retain the feature weights of the clustering identified in the previous step.
11:    end for
12:end for
13:Compute the component-wise median of the retained feature weights.
14:return The indices of the $r$ features with the highest median weights.
4 Experiments setting
We divide our experiments into two main sets to separately evaluate (i) the clustering performance of the proposed MWK++ initialisation method, and (ii) the effectiveness of FS-MWK++ and SFS-MWK++ for unsupervised feature selection.
4.1 Evaluating MWK++
In our first set of experiments, we evaluate whether MWK++ offers improvements over the original Minkowski Weighted $k$-means (MWK), and also compare it against $k$-means++. The latter is the default $k$-means implementation in popular software packages, such as MATLAB, R, and scikit-learn. Hence, it is arguably the de facto clustering algorithm nowadays. We compared these algorithms on synthetic datasets as these offer full control over the data generation process, enabling fair and reproducible comparisons under known ground truth. They also allow us to evaluate performance across a wide range of controlled cluster configurations, which would be unfeasible to obtain with real-world datasets.
We constructed 12 dataset configurations. For each of these configurations we generated 50 datasets, leading to a total of 600 datasets. Each of these datasets contains spherical Gaussian clusters, with each cluster defined by a diagonal covariance matrix whose variance, $\sigma^{2}$, was sampled from a uniform distribution within $[0.5,1.5]$. Hence, we have a mix of dense and sparse clusters. Cluster centroids were independently sampled from a multivariate normal distribution with zero mean and variance of one. The number of data points per cluster was sampled uniformly at random, with a minimum of 20 data points per cluster. To assess robustness under high-dimensional noise, we appended approximately 50% additional noise features to each dataset. These noise features were sampled independently from a uniform distribution.
For example, the dataset configuration 1000x4-5 +2NF contains 50 datasets with 1,000 data points originally described over 4 informative features and partitioned into 5 clusters, with two noise features added, resulting in 6 total features. We evaluated clustering performance using the Adjusted Rand Index (ARI) [24], which quantifies the agreement between predicted and true labels while correcting for chance.
4.2 Evaluating FS-MWK++ and SFS-MWK++
Our second set of experiments focuses on evaluating the effectiveness of our proposed feature selection methods, FS-MWK++ and SFS-MWK++, against well-established unsupervised feature selection baselines. We acknowledge that FS-MWK++ can be computationally expensive for large datasets (this is indeed the reason why we also introduced SFS-MWK++). Hence, we evaluate FS-MWK++ on our synthetic datasets as these are not too large. SFS-MWK++, on the other hand, was designed for scalability. Thus, we evaluate it on much larger real-world datasets we obtained from the popular UCI Machine Learning Repository [25]. We have added approximately 10% and 20% noise features to each real-world dataset. For details, see Table 1.
In the real-world datasets, it is perfectly possible that an original feature of the dataset is not actually relevant. Given we do not have full knowledge of which features are truly relevant (or their degree of relevancy), we also evaluate performance using the Entropy of the resulting clusters (computed using ground-truth labels), under the assumption that lower entropy indicates a purer and more meaningful clustering. We compare our methods against MCFS, SCFS, FSFS, and LSCAE (for details, see Section 2.2).
We normalised all datasets (real-world and synthetic), after adding the noise features, by the range. That is,
$x_{iv}=\frac{x_{iv}-\bar{x}_{v}}{\max(x_{v})-\min(x_{v})},$
where $\bar{x}_{v}$, max($x_{v}$), min($x_{v}$) denote the mean, maximum, and minimum of feature $v$, respectively. We selected the range normalisation instead of the more popular $z$-score because the former does not penalise features with multimodal distributions. This is an important consideration when clusters may be separable along such dimensions.
Table 1: List of the real-world datasets used in our experiments. The column ‘Features’ includes the noise features. We downloaded the datasets from the UCI machine learning repository [25], added uniformly random noise features, and then normalised by the range.
Data points Clusters Features Noise
Dataset $n$ $k$ $m$ Features
CoverType +6NF 581,012 7 60 6
CoverType +11NF 581,012 7 65 11
HandPostures +4NF 78,095 5 40 4
HandPostures +8NF 78,095 5 44 8
IDA2016 +17NF 76,000 2 186 17
IDA2016 +34NF 76,000 2 203 34
OnlineNewsPop +6NF 39,644 6 64 6
OnlineNewsPop +12NF 39,644 6 70 12
SkinSegmentation +1NF 245,057 2 4 1
SkinSegmentation +2NF 245,057 2 5 2
5 Results and discussion
We present our experimental evaluation in two parts. First, we assess the clustering performance of MWK++ against standard baselines. Then, we evaluate the effectiveness of FS-MWK++ and its scalable variant, SFS-MWK++, in unsupervised feature selection tasks.
5.1 Clustering performance of MWK++
Table 2 reports the mean and standard deviation of the Adjusted Rand Index (ARI) across 50 datasets for each configuration. Each algorithm was run 25 times on each dataset, and the reported results reflect the average ARI across all runs. The table compares $k$-means++, MWK (using either the average across all values of the Minkowski exponent $p$, or the best value), and our proposed MWK++ under the same two evaluation schemes.
The results clearly demonstrate that MWK++ consistently outperforms both $k$-means++ and the original MWK across all configurations. Even when comparing MWK++ using all $p$ values (i.e., not selecting the best case), it achieves higher ARI than MWK in all configurations. When using the best exponent $p$, MWK++ yields the highest ARI in every case except one (2000x30-5 +15NF), where $k$-means++ slightly outperforms all variants. This exception is likely due to the relatively small number of clusters, five, combined with a high number of informative features, 30. In this setting, the noise is not sufficient to significantly degrade the performance of $k$-means++.
These results confirm the benefits of our proposed initialisation strategy. The MWK++ method improves centroid selection by leveraging feature relevance estimates early in the clustering process, which leads to more informative separation of clusters. This advantage is especially evident in more challenging scenarios with high dimensionality or large numbers of clusters. For instance, in the configuration 2000x20-20 +10NF, the ARI improves from 0.45 (MWK, best $p$) to 0.66 with MWK++. In addition, the gap between MWK and MWK++ tends to widen as dimensionality and cluster complexity increase. Finally, the relatively low standard deviations across most configurations indicate that MWK++ provides not only better but also stable clustering results compared to its counterparts.
Table 2: Comparison in terms of mean Adjusted Rand Index between the clusterings obtained with $k$-means++, MWK, and MWK++. There are 50 datasets for each configuration. We run each algorithm 25 times on each dataset.
MWK MWK++
$k$means++ All $p$ Best $p$ All $p$ Best $p$
Dataset Mean Std Mean Std Mean Std Mean Std Mean Std
1000x4-3 +2NF 0.02 0.05 0.17 0.06 0.28 0.06 0.31 0.06 0.44 0.05
1000x4-5 +2NF 0.05 0.01 0.15 0.03 0.21 0.03 0.23 0.03 0.32 0.02
1000x4-10 +2NF 0.06 0.00 0.08 0.01 0.11 0.01 0.11 0.01 0.17 0.01
1000x10-3 +5NF 0.20 0.09 0.59 0.05 0.83 0.06 0.71 0.04 0.87 0.06
1000x10-5 +5NF 0.06 0.03 0.42 0.03 0.64 0.02 0.53 0.02 0.70 0.02
1000x10-10 +5NF 0.02 0.01 0.21 0.21 0.33 0.02 0.39 0.01 0.51 0.01
2000x20-5 +10NF 0.54 0.06 0.70 0.02 0.76 0.02 0.85 0.01 0.87 0.01
2000x20-10 +10NF 0.05 0.01 0.53 0.01 0.64 0.01 0.72 0.01 0.78 0.02
2000x20-20 +10NF 0.02 0.01 0.30 0.01 0.45 0.01 0.53 0.01 0.66 0.01
2000x30-5 +15NF 0.94 0.06 0.77 0.02 0.81 0.01 0.87 0.02 0.88 0.03
2000x30-10 +15NF 0.39 0.06 0.62 0.01 0.71 0.01 0.79 0.01 0.83 0.01
2000x30-20 +15NF 0.05 0.01 0.46 0.01 0.61 0.01 0.69 0.02 0.77 0.01
5.2 Unsupervised feature selection
Let us first analyse FS-MWK++. Table 3 presents the results for our FS-MWK++ experiments on synthetic datasets. For each method, we report the average proportion of correctly classified features(those that were either truly informative or correctly identified as noise), along with the corresponding standard deviation. Given these are synthetic datasets we know which features are composed solely of noise.
The results show that FS-MWK++ achieves outstanding performance across all dataset configurations, consistently outperforming MCFS, SCFS, and FSFS, and performing on par with or better than LS-CAE. In fact, FS-MWK++ reaches perfect classification in 8 out of 12 configurations and achieves an average accuracy of 0.99 with a very low standard deviation (0.02), indicating both high precision and stability. This is a strong indication that feature weights derived from MWK++ are highly reliable indicators of feature relevance when aggregated across multiple exponents $p$. While LS-CAE benefits from a learnable deep neural architecture, FS-MWK++ still outperforms it, highlighting the effectiveness of clustering-driven feature weights, even in the absence of supervised learning or end-to-end training. However, LS-CAE did do slightly better in scenarios with only four informative features.
Table 3: Results for the feature selection experiments on synthetic datasets. We present the average proportion of correctly classified features (i.e., informative or non-informative), and related standard deviation.
MCFS SCFS FSFS LSCAE FS-MWK++
Dataset Mean Std Mean Std Mean Std Mean Std Mean Std
1000x4-3 +2NF 0.34 0.05 0.35 0.08 0.94 0.14 1.00 0.00 0.96 0.11
1000x4-5 +2NF 0.34 0.05 0.33 0.00 0.93 0.17 1.00 0.00 0.99 0.07
1000x4-10 +2NF 0.41 0.17 0.33 0.00 0.97 0.12 1.00 0.00 0.98 0.08
1000x10-3 +5NF 0.39 0.09 0.49 0.21 0.88 0.16 0.95 0.07 1.00 0.00
1000x10-5 +5NF 0.36 0.06 0.35 0.06 0.88 0.13 0.92 0.07 1.00 0.00
1000x10-10 +5NF 0.34 0.02 0.33 0.00 0.95 0.08 0.96 0.06 1.00 0.00
2000x20-5 +10NF 0.55 0.11 0.64 0.24 0.89 0.10 0.95 0.04 1.00 0.00
2000x20-10 +10NF 0.38 0.06 0.33 0.00 0.93 0.08 0.96 0.03 1.00 0.00
2000x20-20 +10NF 0.34 0.02 0.33 0.00 0.96 0.05 0.99 0.02 1.00 0.00
2000x30-5 +15NF 0.67 0.07 0.99 0.04 0.89 0.11 0.91 0.04 1.00 0.00
2000x30-10 +15NF 0.62 0.11 0.50 0.24 0.93 0.07 0.96 0.02 1.00 0.00
2000x30-20 +15NF 0.38 0.05 0.33 0.00 0.93 0.08 0.96 0.02 1.00 0.00
Average 0.43 0.07 0.44 0.07 0.92 0.11 0.96 0.03 0.99 0.02
We now turn to the results of our scalable feature selection method, SFS-MWK++, evaluated on real-world datasets with added noise features. We cannot be certain we know which features are informative for each dataset. Hence, we measure performance with the proportion of selected features that were part of the original dataset (as with our previous experiments), and the entropy after feature selection. The latter measures the degree of class purity in each cluster. Lower entropy and higher proportion of correctly identified features both indicate better feature selection. Unfortunately, it was impactical to run SCFS in these large data sets.
SFS-MWK++ performs consistently well across all datasets. It achieves the highest or joint-highest proportion of original features selected in 9 out of 10 datasets, and simultaneously achieves the lowest entropy in the same cases. For example, in CoverType +11NF, SFS-MWK++ selects original features with 98% accuracy and reduces entropy from 2.51 to 1.44, outperforming all competing methods. Similarly strong results are observed on IDA2016 +34NF and OnlineNewsPop +12NF, where SFS-MWK++ both avoids selecting noisy features and yields significantly purer clusters.
These results highlight the effectiveness of SFS-MWK++ in preserving informative features under high-dimensional, noisy conditions. Despite its sampling-based approximation and lack of learning stage or supervision, SFS-MWK++ remains competitive with LS-CAE, which benefits from an end-to-end deep learning framework. Moreover, unlike LS-CAE and FSFS, which show signs of performance degradation in more complex datasets (e.g., IDA2016 +34NF and OnlineNewsPop +12NF), SFS-MWK++ maintains both low entropy and high proportion of original features selected.
Overall, the results in this section support both FS-MWK++ and SFS-MWK++ as effective unsupervised feature selection methods. FS-MWK++ demonstrates near-perfect accuracy in identifying informative and non-informative features in controlled synthetic settings, while SFS-MWK++ extends this success to real-world datasets, achieving strong noise rejection and improved clustering structure with minimal computational overhead. Together, they offer a practical and scalable solution for feature selection in both small and large-scale unsupervised learning tasks.
Table 4: Comparison of feature selection methods on benchmark datasets. For each dataset, we report the original entropy $H_{\text{orig}}$ (computed using ground-truth labels), and for each method, the proportion of selected features that were part of the original dataset, and the entropy after feature selection. A higher proportion suggests better discrimination against artificially added noise features. Lower entropy indicates purer clusters.
FSFS MCFS LS-CAE SFS-MWK++
$H_{\text{orig}}$ Prop. H Prop. H Prop. H Prop. H
CoverType +11NF 2.51 0.91 1.81 0.94 1.54 0.82 2.11 0.98 1.44
CoverType +6NF 2.05 0.93 1.67 0.93 1.53 0.92 1.79 0.97 1.42
HandPostures +4NF 5.31 0.95 5.24 0.90 5.46 1.00 5.02 1.00 5.02
HandPostures +8NF 5.56 0.95 5.24 0.70 5.66 0.95 5.04 1.00 5.02
IDA2016 +17NF 1.94 0.81 2.12 0.96 1.58 0.96 1.49 0.98 1.35
IDA2016 +34NF 2.45 0.66 2.92 0.87 1.83 0.97 1.46 0.97 1.38
OnlineNewsPop +12NF 3.91 0.66 4.50 0.76 3.88 0.80 3.90 0.86 3.54
OnlineNewsPop +6NF 3.53 0.84 3.66 0.91 3.27 0.88 3.58 0.87 3.49
SkinSegmentation +1NF 7.67 0.50 7.70 0.50 7.56 1.00 7.56 1.00 7.56
SkinSegmentation +2NF 7.74 0.20 7.87 0.60 7.56 0.60 7.70 1.00 7.56
6 Conclusion
This paper introduces a novel initialisation method for the Minkowski Weighted $k$-means (MWK) algorithm, we called the Minkowski Weighted $k$-means++ (MWK++). We then used this as a foundational step to design two new unsupervised feature selection algorithms, FS-MWK++ and SFS-MWK++. Our contributions are motivated by the limitations of existing clustering and feature selection methods in high-dimensional, label-free settings, where noise can severely degrade performance. In addition to extensive empirical evaluation, we provide a theoretical guarantee that supports the effectiveness of our feature selection strategy under mild and realistic assumptions.
We showed that MWK++ consistently outperforms both $k$-means++ and the original MWK across a wide variety of synthetic data configurations, especially in high-dimensional or noisy settings. By incorporating feature relevance into the initial centroid selection process, MWK++ improves both clustering accuracy and stability, without introducing significant computational overhead.
Building on this, we proposed FS-MWK++, a feature selection method based on aggregating feature weights across multiple distance exponents. This approach avoids the need for label supervision or deep learning infrastructure while achieving near-perfect discrimination between informative and noisy features in synthetic datasets. To address scalability, we introduced SFS-MWK++, a sampling-based extension that retains the effectiveness of FS-MWK++ while significantly reducing computational cost. Experiments on real-world datasets with added noise confirmed that SFS-MWK++ matches or outperforms state-of-the-art baselines, including the deep learning-based LS-CAE, in terms of both noise rejection and cluster structure.
Overall, our results highlight the potential of clustering-driven feature weighting as a robust foundation for unsupervised learning tasks. Future work will explore theoretical guarantees for weight stability under different $p$ values, and extend our methods to more general clustering frameworks beyond MWK.

### v2

@@
 We support our approach with a theoretical guarantee under mild assumptions and extensive experiments showing that our methods consistently outperform existing alternatives.
+Our software can be found at https://github.com/xzhang4-ops1/FSMWK.
 Keywords: Unsupervised feature selection, clustering, noisy data.
@@
 10:return The indices of the $r$ features with the highest median weights.
-We now formally justify our use of weight stability across different values of $p$ as the basis for FS-MWK++ by presenting a theoretical guarantee. We begin with a short definition and a lemma.
+To establish theoretical guarantees for feature selection, we identify a sufficient condition under which a feature consistently receives high weight across multiple clustering configurations.
+This condition provides the foundation for analysing when FS-MWK++ is guaranteed to select a relevant feature.
 Definition 1.
@@
 Lemma 1.
-Let $v$ be a noise feature, and suppose that all such features are drawn independently from the same distribution and are uncorrelated with cluster structure.
+Let $t$ be a noise feature, and suppose that all such features are drawn independently from the same distribution and are uncorrelated with cluster structure.
 If there exists at least one relevant feature, then for any $p>1$,
-$w_{lv}^{(p)}<\frac{1}{m}.$
+$w_{lt}^{(p)}<\frac{1}{m}.$
 Proof.
-Suppose $u^{*}$ is a relevant feature. Since relevant features are aligned with cluster structure, we have that
-$D_{lu^{*}}^{(p)}<D_{lv}^{(p)}.$
-Let $r=\left(\frac{D_{lv}^{(p)}}{D_{lu^{*}}^{(p)}}\right)^{\frac{1}{p-1}}>1$. Now consider the sum in the denominator of Equation (3),
-$\sum_{u=1}^{m}\left(\frac{D_{lv}^{(p)}}{D_{lu}^{(p)}}\right)^{\frac{1}{p-1}}=r+\sum_{u\neq u^{*}}\left(\frac{D_{lv}^{(p)}}{D_{lu}^{(p)}}\right)^{\frac{1}{p-1}}.$
-Since all noise features have dispersions similar to $D_{lv}^{(p)}$, for each irrelevant $u$,
-$\left(\frac{D_{lv}^{(p)}}{D_{lu}^{(p)}}\right)^{\frac{1}{p-1}}\approx 1\Rightarrow\sum_{u=1}^{m}\left(\frac{D_{lv}^{(p)}}{D_{lu}^{(p)}}\right)^{\frac{1}{p-1}}\approx(m-1)+r>m\Rightarrow w_{lv}^{(p)}<\frac{1}{m}.$
-This lemma allows us to show that, under mild assumptions, relevant features have a higher median weight than noise features when aggregated across clusters and distance exponents.
-The next theorem formalises this insight.
+Let $\mathcal{W}$ be the set of noise features. Then, for a particular cluster $S_{l}$ the denominator of Equation (3) becomes
+$\sum_{u=1}^{m}\left(\frac{D_{lt}^{(p)}}{D_{lu}^{(p)}}\right)^{\frac{1}{p-1}}=\sum_{u\in\mathcal{W}}\left(\frac{D_{lt}^{(p)}}{D_{lu}^{(p)}}\right)^{\frac{1}{p-1}}+\sum_{u\not\in\mathcal{W}}\left(\frac{D_{lt}^{(p)}}{D_{lu}^{(p)}}\right)^{\frac{1}{p-1}}.$
+Given all noise features have similar dispersions,
+$\frac{D_{lt}^{(p)}}{D_{lu}^{(p)}}\approx 1\text{ for all }u\in\mathcal{W}\Rightarrow\sum_{u\in\mathcal{W}}\left(\frac{D_{lt}^{(p)}}{D_{lu}^{(p)}}\right)^{\frac{1}{p-1}}\approx|\mathcal{W}|.$
+Noise features are uncorrelated with cluster structure. Hence, their dispersion will be higher than that of a relevant feature. Thus,
+$\frac{D_{lt}^{(p)}}{D_{lu}^{(p)}}>1\text{ for all }u\not\in\mathcal{W}\Rightarrow\sum_{u\not\in\mathcal{W}}\left(\frac{D_{lt}^{(p)}}{D_{lu}^{(p)}}\right)^{\frac{1}{p-1}}>m-|\mathcal{W}|.$
+Finally,
+$\sum_{u=1}^{m}\left(\frac{D_{lt}^{(p)}}{D_{lu}^{(p)}}\right)^{\frac{1}{p-1}}>m\Rightarrow w_{lt}^{(p)}<\frac{1}{m}.$
+Lemma 2.
+Let $X$ contain both relevant and noise features.
+If each noise feature is drawn independently from a common distribution, then at least one feature will have a weight strictly higher than $\frac{1}{m}$.
+Proof.
+From Lemma 1 we have that if $t$ is a noise feature, then $w_{lt}^{(p)}<\frac{1}{m}$ for any $p>1$ and cluster $S_{l}$.
+Since the feature weights for each cluster $S_{l}$ sum to one, the total deficit $\sum_{t\in\mathcal{W}}\left(\frac{1}{m}-w_{lt}^{(p)}\right)>0$ must be distributed among the remaining (relevant) features.
+Hence, for each pair ($p,l)$ there exists at least one relevant feature $v$ such that $w_{lv}^{(p)}>\frac{1}{m}$.
+∎
+The lemma above guarantees the existence of at least one feature with weight exceeding $\frac{1}{m}$ in some cluster.
+This motivates the next lemma, which characterises the weight of such features at different values of $p$.
+Lemma 3.
+Let $w_{lv}^{(p_{0})}\geq\frac{1}{m}+\gamma$ for some $\gamma>0$. Then $w_{lv}^{(p)}>\frac{1}{m}$ for all $p\in(p_{0}-\delta,\,p_{0}+\delta)$, where
+$\delta=\gamma\cdot\frac{A(p_{0})^{2}(p_{0}-1)^{2}}{L(p_{0})},$
+with
+$A(p_{0})=\sum_{u=1}^{m}\left(\frac{D_{lv}^{(p_{0})}}{D_{lu}^{(p_{0})}}\right)^{1/(p_{0}-1)},\quad L(p_{0})=\sum_{u=1}^{m}\left(\frac{D_{lv}^{(p_{0})}}{D_{lu}^{(p_{0})}}\right)^{\frac{1}{p-1})}\left|\log\left(\frac{D_{lv}^{(p_{0})}}{D_{lu}^{(p_{0})}}\right)\right|.$
+Proof.
+Let $a_{u}=D_{lv}^{(p_{0})}/D_{lu}^{(p_{0})}$. Then
+$w_{lv}^{(p)}=\frac{1}{\sum_{u=1}^{m}\left[\frac{D_{lv}}{D_{lu}}\right]^{\frac{1}{p-1}}}=\frac{1}{\sum_{u=1}^{m}a_{u}^{1/(p-1)}}=\frac{1}{A(p)}.$
+Hence,
+$\frac{d}{dp}w_{lv}^{(p)}=-\frac{A^{\prime}(p)}{A(p)^{2}},\quad\text{where}\quad A^{\prime}(p)=\sum_{u=1}^{m}\frac{a_{u}^{\frac{1}{p-1}}\log a_{u}}{(p-1)^{2}}.$
+Define
+$L(p)=\sum_{u=1}^{m}a_{u}^{\frac{1}{p-1}}|\log a_{u}|,$
+so that
+$|A^{\prime}(p)|\leq\frac{L(p)}{(p-1)^{2}},\quad\text{and}\quad|w_{lv}^{\prime}(p)|\leq\frac{L(p)}{A(p)^{2}(p-1)^{2}}.$
+Let $\delta$ be the minimum perturbation needed to decrease the weight of $v$ to $1/m$. Then by the Mean Value Theorem,
+$\delta=\left(w_{lv}^{(p_{0})}-\frac{1}{m}\right)\cdot\frac{A(p_{0})^{2}(p_{0}-1)^{2}}{L(p_{0})}.$
+Since $w_{lv}^{(p_{0})}\geq\frac{1}{m}+\gamma$,
+$\delta\geq\gamma\cdot\frac{A(p_{0})^{2}(p_{0}-1)^{2}}{L(p_{0})},$
+as claimed. ∎
+Building on these lemmas, we now derive sufficient conditions under which a relevant feature is guaranteed to be selected by FS-MWK++.
+These conditions link the smoothness of feature weights across $p$, the proportion of clusters where the feature appears prominent, and the structure of the $P$ grid.
+The result formalises the intuitive idea that stability and alignment with cluster structure make a feature likely to stand out.
 Theorem 1.
-Let $P\neq\emptyset$ be a finite set of valid Minkowski exponents.
-If $u^{*}$ is a relevant feature for at least half of the clusters and $v$ is a noise feature drawn independently from a common distribution and uncorrelated with cluster structure, then
-$\mathrm{median}\left(\{w_{lu^{*}}^{(p)}:p\in P,\,l=1,\dots,k\}\right)>\mathrm{median}\left(\{w_{lv}^{(p)}:p\in P,\,l=1,\dots,k\}\right).$
+Let $X$ contain both relevant and noise features, and suppose noise features are drawn independently from a common distribution and are uncorrelated with cluster structure.
+Let $v$ be a feature such that for a proportion $\alpha$ of the $k$ clusters, there exists a $p_{l}\in P$ satisfying
+$w_{lv}^{(p_{l})}\geq\frac{1}{m}+\gamma,$
+with $\gamma>0$ and suppose that for all such $p_{l}$,
+$\gamma\cdot\frac{A(p_{l})^{2}(p_{l}-1)^{2}}{L(p_{l})}>\frac{1}{2\alpha}.$
+Then FS-MWK++ is guaranteed to select $v$, provided that the number of selected features $r$ is at least the number of features satisfying this condition.
 Proof.
-From Lemma 1, we have that for any $p>1$
-$w_{lv}^{(p)}\in\{w_{lv}^{(p)}:p\in P,\,l=1,\dots,k\}\Rightarrow w_{lv}^{(p)}<\frac{1}{m}.$
-Hence, $\mathrm{median}\left(\{w_{lv}^{(p)}:p\in P,\,l=1,\dots,k\}\right)<\frac{1}{m}$.
-Now consider the set $\{w_{lu^{*}}^{(p)}:p\in P,\,l=1,\dots,k\}$. Observe that Equation (3) is strictly convex for $p>1$, so
-$w_{lu^{*}}^{(p)}=\frac{1}{m}\iff\sum_{u=1}^{m}\left[\frac{D_{lu^{*}}}{D_{lu}}\right]^{\frac{1}{p-1}}=m\iff\frac{D_{lu^{*}}}{D_{lu}}=1\text{ for all }u=1,\ldots,m.$
-However, the existence of $v$ shows $X$ has at least one noise feature.
-Thus, the above cannot happen and by consequence $w_{lu^{*}}^{(p)}\neq\frac{1}{m}$.
-Recall that (3) ensures the weights for a particular cluster $S_{l}$ must sum to one, and that $w_{lv}^{(p)}<\frac{1}{m}$.
-Hence, the difference $\frac{1}{m}-w_{lv}^{(p)}>0$ is spread over any relevant feature, and $w_{lu^{*}}^{(p)}>\frac{1}{m}$.
-We have, by hypothesis, that $u^{*}$ is relevant for at least half of the clusters.
-Therefore, $\mathrm{median}\left(\{w_{lu^{*}}^{(p)}:p\in P,\,l=1,\dots,k\}\right)>\frac{1}{m}$.
-∎
+Notice there are $|P|\cdot k$ pairs $(p,l)$.
+To guarantee $v$ is selected, it is sufficient to show that $w_{lv}^{(p_{l})}>\frac{1}{m}+\gamma$ for some $\gamma>0$ in at least $\frac{|P|\cdot k}{2}$ pairs.
+Let this weight condition hold for $\alpha\cdot k$ clusters for each $p_{l}$ in $\mathcal{P}$.
+Then,
+$\alpha\cdot k\cdot|\mathcal{P}|>\frac{|P|\cdot k}{2}\iff 2\alpha|\mathcal{P}|>|P|\iff|\mathcal{P}|>\frac{|P|}{2\alpha}.$
+Recall that $|P|=20$ and that the interval $[p-\delta,p+\delta]$ has length $2\delta$. Given the step within $P$ is of $0.1$, we have
+$|\mathcal{P}|\approx\frac{2\delta}{0.1}=20\delta=\delta\cdot|P|>\frac{|P|}{2\alpha}\iff\delta>\frac{1}{2\alpha}.$
+Following Lemma 3, the above leads to
+$\gamma\cdot\frac{A(p_{l})^{2}(p_{l}-1)^{2}}{L(p_{l})}>\frac{1}{2\alpha}.$
+Let us now demonstrate Theorem 1 with an example.
+Let $m=12$, $\gamma=0.15$, and $\alpha=0.9$.
+That is, there exists a feature $v$ such that for 90% of the clusters $S_{l}\in S$ there exists a (possibly different) $p_{l}\in P$ for which $w_{lv}^{(p_{l})}\geq\frac{1}{m}+\gamma$.
+Recall that FS-MWK++ considers a total of $|P|\cdot k$ pairs $(p,l)$.
+In this example, the feature is stated to have a high weight in only $0.9\cdot k$ pairs (as there is one $p_{l}$ for each cluster).
+Hence, it is not immediately obvious FS-MWK++ would select $v$ under mild assumptions.
+Notice that
+$w_{lv}^{(p_{l})}=\frac{1}{A(p_{l})}\geq\frac{1}{m}+\gamma\Rightarrow A(p_{l})\leq\frac{1}{\frac{1}{m}+\gamma}=\frac{1}{\frac{1}{12}+0.15}\approx 4.29.$
+Let us assume a value of approximately 80%, $A(p_{l})=3.4$, and $p_{l}=2.4$. Recall that $a_{u}=D_{lv}^{(p_{l})}/D_{lu}^{(p_{l})}$, then
+$A(p_{l})=\sum_{u=1}^{m}a_{u}^{1/(p_{l}-1)}=\sum_{u=1}^{12}a_{u}^{1/1.4}=3.4,$
+leading to, on average, $a_{u}^{1/1.4}\approx 0.283\iff a_{u}\approx 0.171$. Let us use this to approximate $L(p_{l})$,
+$L(p_{l})=\sum_{u=1}^{m}a_{u}^{\frac{1}{p-1}}|\log a_{u}|\approx\sum_{u=1}^{12}0.283|\log 0.171|\approx 6.00.$
+Then,
+$\gamma\cdot\frac{A(p_{l})^{2}(p_{l}-1)^{2}}{L(p_{l})}>\frac{1}{2\alpha}\iff 0.15\frac{3.4^{2}\cdot(2.4-1)^{2}}{6.00}\approx 0.566>\frac{1}{2}.$
+Therefore, the condition in Theorem 1 is satisfied, and $v$ is guaranteed to be selected by FS-MWK++.
 While FS-MWK++ is effective at identifying stable and informative features (for details, see Section 5), it becomes computationally expensive on large datasets.
@@
 We divide our experiments into two main sets to separately evaluate (i) the clustering performance of the proposed MWK++ initialisation method, and (ii) the effectiveness of FS-MWK++ and SFS-MWK++ for unsupervised feature selection.
+The source code for our experiments can be found at https://github.com/xzhang4-ops1/FSMWK.
 4.1 Evaluating MWK++
@@
 Overall, our results highlight the potential of clustering-driven feature weighting as a robust foundation for unsupervised learning tasks.
-Future work will explore theoretical guarantees for weight stability under different $p$ values, and extend our methods to more general clustering frameworks beyond MWK.
+Future work will explore further theoretical guarantees, and extend our methods to more general clustering frameworks beyond MWK++.

### v3

@@
 Clustering is a fundamental machine learning technique that assigns data points in a dataset to clusters, with each cluster containing similar data points.
-As clustering algorithms operate without the need for labelled data, they have been successfully applied across various domains, including data pre-processing, quantitative finance, image analysis, and bioinformatics [1, 2, 3, 4].
+As clustering algorithms operate without the need for labelled data, they have been successfully applied across various domains, including data pre-processing, quantitative finance, image analysis, and bioinformatics [1, 2, 3, 4, 5].
 However, their performance (and that of any other machine learning algorithm) can be heavily influenced by the quality and dimensionality of the data.
@@
 2.1 Clustering
-The $k$-means algorithm [9] is a widely used partitional clustering method, which has been extended in numerous ways [13, 11].
+The $k$-means algorithm [10] is a widely used partitional clustering method, which has been used and extended in numerous ways [14, 15, 16, 12].
 It produces a partition $S=\{S_{1},\ldots,S_{k}\}$ of a dataset $X$ such that $S_{l}\cap S_{t}=\emptyset$ for all $l,t=1,\ldots k$ with $l\neq t$.
@@
 It improves the initial centroids of $k$-means by spreading them out more strategically based on the data distribution.
-This often reduces the likelihood of poor local minima, leading to better clustering results.
+This often reduces the likelihood of poor local minima, leading to better clustering results [18, 19].
 The $k$-means++ algorithm is now the default $k$-means implementation in many popular software packages such as MATLAB, scikit-learn, and R.
@@
 Feature selection addresses these challenges by identifying such features, thereby leading to more efficient, stable, and reliable results.
-Feature selection using feature similarity (FSFS) [20] is one of the most popular unsupervised feature selection methods (as indicated by its high number of citations in Google Scholar).
+Feature selection using feature similarity (FSFS) [25] is one of the most popular unsupervised feature selection methods [26, 27, 28], which is also indicated by its high number of citations in Google Scholar.
 Hence, we include it in our comparison (see Section 5).
@@
 Here, this use of $\kappa$ is unrelated to the $k$ of $k$-means.
-Multi-Cluster Feature Selection (MCFS) [21] is another popular unsupervised feature selection method.
+Multi-Cluster Feature Selection (MCFS) [29] is another popular unsupervised feature selection method [6, 30].
 MCFS combines spectral techniques from manifold learning with $\ell_{1}$-regularization to identify the most relevant features.
@@
 Subspace Clustering Feature Selection (SCFS) [31] is a recent unsupervised feature selection method combining subspace clustering with sparse regression.
-SCFS aims to select features that preserve the multi-cluster structure of the data by adaptively learning sample similarities.
+SCFS aims to select features that preserve the multi-cluster structure of the data by adaptively learning sample similarities [32].
 First, it computes a similarity matrix implicitly by solving a self-expressive model, in which each sample is represented through a low-dimensional space shared with other similar samples.
@@
 We direct readers interested in further details about SCFS to its original publication [31].
-Laplacian Score-regularized Concrete Autoencoder (LS-CAE) [23] is a recent deep learning-based approach to unsupervised feature selection.
+Laplacian Score-regularized Concrete Autoencoder (LS-CAE) [33] is a recent deep learning-based approach to unsupervised feature selection [26].
 LS-CAE extends the Concrete Autoencoder (CAE) framework by incorporating a Laplacian score regularisation term into the loss function.